# A multi-stage Bayesian approach to fit spatial point process models

### v1

@@ 5.1 Neural Network Basis Expansion
where $g(\cdot)$ represents a user-specified activation function, $\boldsymbol{\omega}_{j}$ is a $p\times 1$ vector of weights, and $\tilde{\beta}_{j}$ represents the coefficient associated with the $j$th hidden node. We omitted a bias term in (20) because we found that the weights were sufficient for constructing the basis expansion. The quantity in (20) can be summarized compactly as
@@
The matrix $\mathbf{W}$ is referred to as the hidden layer output matrix. Instead of learning the weights, $\boldsymbol{\omega}_{j}$ for $j=1,\dots,q$ using gradient-based methods, the ELM scheme samples random weights from a user-specified distribution for efficient computation Huang et al. (2006). After we computed $\mathbf{W}$ using the randomly generated weights, we used it as a basis expansion for $\mathbf{X}$ (i.e., each hidden node represents a basis vector), which then served as a new design matrix for the first-stage logistic regression problem. Aside from replacing $\mathbf{X}$ with $\mathbf{W}$, the intermediate and second stages in the multi-stage algorithm remain unchanged.
For the 21 June 2007 harbor seal pup data, we used the Gaussian Error Linear Units (GELU) activation function (Hendrycks and Gimpel, 2016) defined as $g(\mathbf{x})=\mathbf{x}\Phi(\mathbf{x})$, where $\Phi(x)$ is a standard Gaussian cumulative distribution function, and generated 100 hidden layer output matrices for $q=5$ and weights sampled from a standard Gaussian distribution. We selected the $\mathbf{W}$ matrix with the optimal AIC based on (20) for the basis expansion. We then used the GLM-E multi-stage algorithm to obtain posterior realizations for $\tilde{\boldsymbol{\beta}}$, which were used to simulate points across the entire study domain using the method described in section 3.3.2. The simulated points were summarized in a grid of posterior mean abundance (Figure 6). We validated our model using posterior predictive model-checking and did not observe large discrepancies between the simulated and observed data (Appendix D).
@@
When compared to the observed point pattern in Figure 5, Figure 6 indicates that greater abundance values were simulated in areas of high observed pup concentrations in addition to surrounding areas that were originally unobserved. Areas of greater simulated abundance values are also consistent with areas of higher ice proportions and more negative bathymetry values (i.e., areas of deeper water), supporting previous studies on harbor seal behavior in Southeast Alaska (Montgomery et al., 2007; Kaluzienski et al., 2023). Additionally, although we did not include this as a covariate, greater abundance values were simulated in areas where eddies typically form in the inlet, which is consistent with findings that harbor seals prefer relatively slow-moving icebergs during the pupping season (Kaluzienski et al., 2025). The posterior predictive distribution for total abundance $N$ was generated using the procedure in 3.3.1 (Figure 7). The posterior mean and standard deviation of NN were 835.2 and 27.7, respectively.
@@ 6 Discussion
Recent innovation in technology and survey techniques have led to significant improvements in the availability and spatial resolution of ecological datasets, making the development of computationally-efficient and scalable spatial point process models especially crucial. Many useful methods have been developed to facilitate flexible and efficient SPP modeling (Baddeley et al., 2010; Warton and Shepherd, 2010; Aarts et al., 2012; Fithian and Hastie, 2013) for conditional IPP likelihoods as in (2); however, fitting a Bayesian IPP model using the complete likelihood in (1), which accounts for the stochasticity associated with $n$, using conventional MCMC algorithms has remained computationally intensive.
@@
More recently, recursive Bayes methods for fitting ecological statistical models have gained traction and have been shown to substantially improve computational efficiency (Hooten et al., 2021, 2023, 2024; Leach et al., 2022; McCaslin et al., 2021). We showed that the complete likelihood for an IPP with spatial covariates can be similarly fit in stages where the coefficients are first estimated using the likelihood conditioned on $n$, then corrected with the Poisson likelihood for $n$.
@@
Fitting the IPP model in stages provides many benefits including increased flexibility in the first stage. We showed that fitting the first stage simplifies to fitting a model using the conditional likelihood (2), which is well approximated using logistic regression. This approximation provides various ways to fit the first stage, including user-friendly GLM R packages, making fitting Bayesian SPP models more accessible to a broader audience. Additionally, the multi-stage methods we presented do not require algorithm supervision and result in improved mixing for MCMC chains when compared to a single-stage MCMC algorithm, especially for the non-Bayesian first-stage approaches, GLM-E and GLM-A. Furthermore, all aforementioned multi-stage methods leverage parallel computing resources to pre-compute computationally intensive values, allowing the method to scale with the number of computer cores and become more feasible to fit for a large number of quadrature points.
The fast non-Bayesian estimation for logistic regression also allows for convenient pre-training of NN basis functions using the ELM scheme. A similar strategy can be applied to training other basis functions, which can be found in Hefley et al. (2017). Furthermore, tuning the parameters for a single-stage random walk MCMC algorithm becomes cumbersome as the number of basis vectors increases. Thus, it is especially useful to implement unsupervised MCMC algorithms, such as those that utilize Gibbs sampling and/or optimization-based estimation, when using a basis expansion.
@@
One main limitation of modeling the points as an IPP is that points are assumed conditionally independent given the spatial covariates. Basis functions can allow for complex and highly non-linear relationships between covariates and point intensity, but additional dependence among points may be necessary to reasonably capture some extreme forms of spatial heterogeneity such as clustering due to social structure. Future directions of this work include extending the multi-stage procedure for self-exciting point processes such as the Hawkes process (Hawkes, 1971).
@@ 7 Acknowledgments
Recall that the full likelihood is given by
@@
We can rewrite the full likelihood using $\zeta:=e^{\beta_{0}}$:
@@
Recall that $\Lambda(\mathcal{S})$ in the full-conditional update can be pre-computed in the previous stage in the multi-stage MCMC algorithm. Thus, sampling from the full-conditional distribution for $\zeta$ is fast.
@@
We used three covariates to fit the IPP to harbor seal pup data on 21 June 2007: bathymetry, ice proportion, and distance to glacier. For the bathymetry covariate, we downloaded a raster with cell size 1/3 arc-second ($\approx$ 10 m) from the National Oceanic and Atmospheric Administration (NOAA) Hydrographic Survey Metadata Database (HSMD). The raster values are relative to Mean Lower Low Water, with more negative values corresponding to deeper waters. We used the bathymetry raster grid to construct the remaining covariate rasters. Ice data were collected using the same aerial imagery survey described in section 5. Kaluzienski et al. (2023) conducted an ice segmentation approach to obtain iceberg outlines within each image footprint. We used the iceberg outlines to compute iceberg proportions and performed kriging to complete the ice proportion raster (Appendix C). To construct the the distance to glacier terminus raster, we computed the Euclidean distance from each grid center to the closest point on the glacier terminus.
@@
The goals for utilizing the aerial ice imagery were to (1) quantify ice availability from the raw images by creating a rasterized map of ice proportions and (2) produce a map that has the same resolution and spatial extent as the other covariates being used in the model (i.e., bathymetry and distance from glacier terminus) across the entire study domain. To ensure an appropriate grid to account for the varying locations of observed ice proportions, the geometries of each image footprint in the survey date were first intersected with a pre-specified raster (Figure 8A). The resulting output was a series of grid cell geometries and their corresponding centroid coordinates.
Using these grid cell geometries, as well as a binarized version of the observed icebergs, an iterative masking method was used to find the proportion of ice within each grid cell. For each image in the survey date, the raster grid was superimposed onto the image and a logical mask was used to iterate through the grid cell and image intersections. Within a grid cell, the total number of iceberg pixels was divided by the total number of pixels in the grid cell, resulting in an ice proportion value. This process was repeated for each grid cell in an image, and for each image in the survey date (Figure 8B). Because ice was only observed within the areas of the image footprints, the calculated ice proportions and the LatticeKrig R package were used to produce an interpolated surface of ice proportions across the entire survey domain (Nychka et al., 2023). After performing kriging, a raster was produced, where ice proportion estimates were available at the same resolution as the other covariates used, and across the entire survey domain.
@@
Appendix D
To check whether the observed data reasonably arises from the estimated inhomogeneous Poisson process, we performed a posterior predictive check based on the L-function, a function used to summarize the level of clustering in a spatial point pattern at varying distances. For each $\beta_{0}^{(k)}$ and $\boldsymbol{\beta}^{(k)}$ posterior realization obtained from the multi-stage MCMC method, we simulated points within the image footprints and computed the corresponding L-function using an isotropic edge correction. These were then compared to the L-function computed for the observed harbor seal pup locations. We found that the L-function computed from the observed pup data lie within the realm of posterior predictive L-functions (Figure 9).
Figure 9: Posterior predictive L-functions (gray) compared to the observed L-function (red).

### v2

@@ 5.1.1 Neural Network Basis Expansion
where $g(\cdot)$ represents a user-specified activation function, $\boldsymbol{\omega}_{j}$ is a $p\times 1$ vector of weights, and $\tilde{\beta}_{j}$ represents the coefficient associated with the $j$th hidden node. The quantity in (20) can be summarized compactly as
@@
The matrix $\mathbf{W}$ is referred to as the hidden layer output matrix. Instead of learning the weights, $\boldsymbol{\omega}_{j}$ for $j=1,\dots,q$ using gradient-based methods, the ELM scheme samples random weights from a user-specified distribution for efficient computation (Huang et al., 2006). After we computed $\mathbf{W}$ using the randomly generated weights, we used it as a basis expansion for $\mathbf{X}$ (i.e., each hidden node represents a basis vector), which then served as a new design matrix for the first-stage logistic regression problem. Aside from replacing $\mathbf{X}$ with $\mathbf{W}$, the intermediate and second stages in the multi-stage algorithm remain unchanged.
For the 21 June 2007 harbor seal pup data, we used the Gaussian Error Linear Units (GELU) activation function (Hendrycks and Gimpel, 2016) defined as $g(\mathbf{x})=\mathbf{x}\Phi(\mathbf{x})$, where $\Phi(x)$ is a standard Gaussian cumulative distribution function, and generated 100 hidden layer output matrices for $q=5$ and weights sampled from a standard Gaussian distribution. We selected the $\mathbf{W}$ matrix with the optimal AIC based on (20) for the basis expansion. We then used the GLM-E multi-stage algorithm to obtain $K=100,000$ posterior realizations for $\tilde{\boldsymbol{\beta}}$ on the same machine as the simulation study (24-core, 3.68 GHz processors and 192 GB of RAM), which required approximately three minutes for the entire multi-stage algorithm. The posterior realizations were used to simulate points across the entire study domain using the method described in section 3.3.2. The simulated points were summarized in a grid of posterior mean abundance (Figure 8). We validated our model using posterior predictive model-checking and did not observe large discrepancies between the simulated and observed data (Appendix E).
@@
When compared to the observed point pattern in Figure 7, Figure 8 indicates that greater abundance values were simulated in areas of high observed pup concentrations in addition to surrounding areas that were originally unobserved. Areas of greater simulated abundance values are also consistent with areas of higher ice proportions and more negative bathymetry values (i.e., areas of deeper water), supporting previous studies on harbor seal behavior in Southeast Alaska (Montgomery et al., 2007; Kaluzienski et al., 2023). Additionally, although we did not include this as a covariate, greater abundance values were simulated in areas where eddies typically form in the inlet, which is consistent with findings that harbor seals prefer relatively slow-moving icebergs during the pupping season (Kaluzienski et al., 2025). The posterior predictive distribution for total abundance $N$ was generated using the procedure in 3.3.1 (Figure 9). The posterior mean and standard deviation of $N$ were 835.2 and 27.7, respectively.
5.1.2 Alternative Strategies for Increased Model Flexibility
Although we found the NN basis expansion approach to be both sufficient for capturing complex spatial patterns, such as the harbor seal pup data, and computationally efficient within the multi-stage computing algorithm, alternative strategies for increased model flexibility exist and may be applicable. For instance, fixed-rank or radial basis functions may be used instead of the NN basis functions. Dovers et al. (2024) showed that a similar basis function approach can be used to approximate an LGCP model and only requires fitting a generalized linear mixed model (GLMM) in R. Fitting the multi-stage algorithm using an SPP model with a true latent Gaussian process to capture additional spatial heterogeneity as in an LGCP model (Brix and Diggle, 2001) is achievable, but introduces additional computational complexities associated with learning parameters for a high-dimensional latent field, which is the subject of ongoing research.
@@ 6 Discussion
Recent innovation in technology and survey techniques have led to significant improvements in the availability and spatial resolution of datasets, making the development of computationally-efficient and scalable SPP models especially crucial. Many useful methods have been developed to facilitate flexible and efficient SPP modeling (Baddeley et al., 2010; Warton and Shepherd, 2010; Aarts et al., 2012; Fithian and Hastie, 2013) for conditional IPP likelihoods as in (2). Furthermore, developments in using the Laplace approximation to approximate Bayesian SPP model marginal distributions have led to significant computational improvements (Rue and Martino, 2009; Illian et al., 2012) for fitting LGCP models. However, fitting an SPP model that both uses the complete likelihood (1), which accounts for the stochasticity associated with $n$, and provides exact Bayesian inference for the joint posterior distribution has remained computationally challenging.
@@
More recently, recursive Bayes methods for fitting ecological statistical models have gained traction and have been shown to substantially improve computational efficiency (Hooten et al., 2021, 2023, 2024; Leach et al., 2022; McCaslin et al., 2021). We showed that the complete likelihood for an IPP model (1) can be similarly fit in stages where the coefficients are first estimated using the likelihood conditioned on $n$ (2), then corrected with the Poisson likelihood for $n$ (8). Although sample degeneracy is a commonly cited issue for recursive Bayes and other sequential filtering algorithms (Barreto and Hooten, 2025; Scharf, 2025; Tyalor et al., 2025), our multi-stage algorithm did not suffer from this phenomenon due to only having two stages and the ability to sample the additional learned parameter in the second stage, $\beta_{0}$, using a Gibbs sampler. This is especially true for the non-Bayesian first stage strategies, GLM-A and GLM-E, because the transient posterior realizations from (16) are independent. In practice, the acceptance rate for $\boldsymbol{\beta}$ in the second stage ranged from 0.25 to 0.7 in our simulations and case study.
@@
Fitting the IPP model in stages provides many benefits including increased flexibility in the first stage. We showed that fitting the first stage is approximately equivalent to fitting a model using the conditional likelihood (2), which is well approximated using logistic regression. This approximation provides various ways to fit the first stage, including user-friendly GLM R packages, making fitting Bayesian SPP models more accessible to a broader audience. Additionally, the multi-stage methods we presented do not require algorithm supervision and result in improved mixing for MCMC chains when compared to a single-stage MCMC algorithm, especially for the non-Bayesian first-stage approaches, GLM-E and GLM-A. Furthermore, all aforementioned multi-stage methods leverage parallel computing resources to pre-compute computationally intensive values, allowing the method to scale with the number of computer cores and increase the feasibility of fitting an IPP model using a large number of quadrature points.
The multi-stage construction of the algorithm also allows practitioners to take a meta-analytic approach to adjust model output that was originally fit using the conditional IPP likelihood (2) by incorporating data that are already available. The original MCMC sample (or approximate Gaussian distribution if non-Bayesian) can be treated as the transient posterior realizations in the multi-stage algorithm. Proceeding with the intermediate and second stages adjusts the estimates to account for the stochasticity associated with $n$ and provides learning about the $\beta_{0}$ parameter, which was non-identifiable during the original model fitting. The resulting model output provides exact Bayesian inference using the complete likelihood in (1).
@@
Restricting the IPP intensity function to be log-linear in the available covariates may not be sufficient to capture spatial heterogeneity in complex spatial patterns, such as the harbor seal pup data in Figure 6. In section 5.1, we outlined ways to account for additional spatial heterogeneity in the model, including using a NN basis expansion, for which training is compatible and efficient within the multi-stage algorithm. Furthermore, adding x- and y-coordinates to the basis expansion functions similarly to the LGCP in that they account for spatial heterogeneity that is unexplained by the observed covariates Dovers et al. (2024). One main limitation of modeling the points as an IPP or LGCP is that points are assumed conditionally independent given the spatial covariates (and latent field for the LGCP). Basis functions and/or latent fields can allow for flexible estimation of complex intensity fields, but additional dependence among points may be necessary to reasonably capture extreme forms of spatial heterogeneity such as clustering due to social structure. Exploratory data analysis on the observed point pattern and model validation techniques concerned with the second-order properties of a point pattern (Illian et al., 2008) should be explored during analysis. Future directions of this work include extending the multi-stage procedure for self-exciting point processes such as the Hawkes process (Hawkes, 1971).
@@ 7 Acknowledgments
Recall that the complete likelihood is given by
@@
We can rewrite the complete likelihood using $\zeta:=e^{\beta_{0}}$:
@@
Recall that $\Lambda(\mathcal{S})$ in the full-conditional update can be pre-computed in the previous stage in the multi-stage MCMC algorithm. Thus, sampling from the full-conditional distribution for $\zeta$ in the second stage is fast.
@@
We conducted simulations using two distinct compact windowed constructions in which the windows cover approximately half of the study domain area. For each of the windowed constructions, we fit the IPP model with the windowed complete likelihood (3) using a single-stage MCMC algorithm and each multi-stage method to data generated from two distinct latent intensity fields. The simulation study with small windows and point pattern 1 is presented in section 4, while the remaining three simulations are shown below.
Small windows, point pattern 2
Figure 10: (a) Intensity heat map for simulated data. (b) Realization of simulated points. Points located within the compact windows (red; $n$ = 617) were used to fit the model.
Figure 12: Comparison of bivariate joint posterior distributions for each pair of model parameters. Contours represent kernel density estimates of the joint posterior distributions. The GLM-E contours (pink) are overlaid on top of the single-stage contours (gray).
$\beta_{0}$ $\beta_{1}$ $\beta_{2}$
Single-stage 0.170 0.167 0.146
PG 0.038 0.032 0.045
HMC 0.017 0.015 0.020
GLM-E 0.001 0.002 0.002
GLM-A 0.001 0.001 0.002
Large windows, point pattern 1
Figure 14: (a) Intensity heat map for simulated data. (b) Realization of simulated points. Points located within the compact windows (red; $n$ = 528) were used to fit the model.
Figure 16: Comparison of bivariate joint posterior distributions for each pair of model parameters. Contours represent kernel density estimates of the joint posterior distributions. The GLM-E contours (pink) are overlaid on top of the single-stage contours (gray).
$\beta_{0}$ $\beta_{1}$ $\beta_{2}$
Single-stage 0.128 0.139 0.137
PG 0.028 0.029 0.035
HMC 0.015 0.015 0.018
GLM-E 0.001 0.001 0.001
GLM-A 0.001 0.001 0.001
Large windows, point pattern 2
Figure 18: (a) Intensity heat map for simulated data. (b) Realization of simulated points. Points located within the compact windows (red; $n$ = 483) were used to fit the model.
Figure 20: Comparison of bivariate joint posterior distributions for each pair of model parameters. Contours represent kernel density estimates of the joint posterior distributions. The GLM-E contours (pink) are overlaid on top of the single-stage contours (gray).
$\beta_{0}$ $\beta_{1}$ $\beta_{2}$
Single-stage 0.162 0.174 0.157
PG 0.040 0.037 0.048
HMC 0.023 0.021 0.027
GLM-E 0.002 0.002 0.002
GLM-A 0.001 0.001 0.001
@@
We used three covariates (bathymetry, ice proportion, and distance to glacier) to fit the IPP model to harbor seal pup data on 21 June 2007. For the bathymetry covariate, we downloaded a raster with cell size 1/3 arc-second ($\approx$ 10 m) from the National Oceanic and Atmospheric Administration (NOAA) Hydrographic Survey Metadata Database (HSMD). The raster values are relative to Mean Lower Low Water, with more negative values corresponding to deeper waters. We used the bathymetry raster grid to construct the remaining covariate rasters. Ice data were collected using the same aerial imagery survey described in section 5. Kaluzienski et al. (2023) conducted an ice segmentation approach to obtain iceberg outlines within each image footprint. We used the iceberg outlines to compute iceberg proportions and performed kriging to complete the ice proportion raster (Appendix D). To construct the the distance to glacier terminus raster, we computed the Euclidean distance from each grid center to the closest point on the glacier terminus.
Appendix D
The goals for utilizing the aerial ice imagery were to (1) quantify ice availability from the raw images by creating a rasterized map of ice proportions and (2) produce a map that has the same resolution and spatial extent as the other covariates used in the model (i.e., bathymetry and distance from glacier terminus) across the entire study domain. To ensure an appropriate grid to account for the varying locations of observed ice proportions, the geometries of each image footprint in the survey date were first intersected with a pre-specified raster (Figure 9A). The resulting output was a series of grid cell geometries and their corresponding centroid coordinates.
Using these grid cell geometries, as well as a binarized version of the observed icebergs, an iterative masking method was used to find the proportion of ice within each grid cell. For each image in the survey date, the raster grid was superimposed onto the image and a logical mask was used to iterate through the grid cell and image intersections. Within a grid cell, the total number of iceberg pixels was divided by the total number of pixels in the grid cell, resulting in an ice proportion value. This process was repeated for each grid cell in an image, and for each image in the survey date (Figure 22B). Because ice was only observed within the areas of the image footprints, the calculated ice proportions and the LatticeKrig R package (Nychka et al., 2023) were used to produce an interpolated surface of ice proportions across the entire survey domain. After performing kriging, a raster was produced, where ice proportion estimates were available at the same resolution as the other covariates used, and across the entire survey domain.
@@
Appendix E
To check whether the observed data reasonably arises from the fitted IPP model, we performed a posterior predictive check based on the L-function, a function used to summarize the level of clustering in a spatial point pattern at varying distances. For each $\beta_{0}^{(k)}$ and $\boldsymbol{\beta}^{(k)}$ posterior realization obtained from the multi-stage GLM-E method, we simulated points within the image footprints and computed the corresponding L-function using an isotropic edge correction. These were then compared to the L-function computed for the observed harbor seal pup locations.
We performed this model checking technique for the model with and without the NN basis expansion. For the model without the NN basis expansion, we found that the L-function computed from the observed pup data did not lie within the realm of posterior predictive L-functions (Figure 23), indicating that the model did not sufficiently account for clustering in the observed point pattern. However, after refitting the model with the inclusion of the NN basis expansion, we found that the L-function computed from the observed pup data lie within the realm of posterior predictive L-functions (Figure 24), indicating a reasonable model fit.
Figure 23: Posterior predictive L-functions from the model without NN basis expansion (gray) compared to the observed L-function (red).
Figure 24: Posterior predictive L-functions from the model with NN basis expansion (gray) compared to the observed L-function (red).